OpenStudy (thedj4jc):

Describe the translation of the graph of y = x^2 that results in the graph of y = (x - 3)^2. left 3 units right 3 units down 3 units up 3 units

OpenStudy (thedj4jc):

I think it's C. down 3 units. Am I right?

OpenStudy (phi):

you shift up or down by adding a number (call it k), like this \[ y = x^2 + k\] to move the graph sideways you shift by h this way \[ y = (x-h)^2 \] the trick part is which way , left or right?

OpenStudy (thedj4jc):

It'd be left, I think. Because the negative numbers are on the left, and the minus/negative sign would shift it left? I think.......

OpenStudy (thedj4jc):

@phi

OpenStudy (phi):

sorry for the long delay. For the y shift (up and down) it works they way you think. +k means shift up k steps and -k means down k steps but for the sideways shift , it works *backwards* (gack!) so y=(x-3)^2 shifts the graph y= x^2 three steps to the *right*